It takes Eddie 7 hours to paint a wall. Alison takes 4 hurs to do the same job. how long would it take them, working together, to paint the wall? Need breakdown please.
Found 2 solutions by scott8148, Alan3354:
Answer by scott8148(6628)

in the x hrs it takes for them to paint the wall, Eddie does x/7 of the job and Alison does x/4
___ these two fractions sum to one, the whole job

(x/7) + (x/4) = 1

multiplying by LCD (28) ___ 4x + 7x = 28 ___ 11x = 28 ___ x = 28/11

Answer by Alan3354(30993)

Use product/sum
= 7*4/(7+4) = 28/11 hours